OpenStudy (anonymous):

What is the molarity of a NaCL solution that has 58 grams of NaCl dissolved in 4.0 liters of water?

OpenStudy (dtan5457):

molarity= # of moles/liters of solute

OpenStudy (anonymous):

How many moles is it if there is 58 grams? (Or how do I find that out?)

OpenStudy (dtan5457):

Find the molar mass of NaCl. given mass/formula mass

OpenStudy (dtan5457):

Na=roughly 22 grams Na=roughly 35 grams

OpenStudy (dtan5457):

*cl is roughly 35 grams

OpenStudy (anonymous):

So 58/57=1 ?

OpenStudy (dtan5457):

Yeah, basically.

OpenStudy (dtan5457):

Now, there is 4 liters in the solute. 1/4=? from there, just remember molarity is measured as M

OpenStudy (thadyoung):

First you need to get the grams given into moles, because molarity is defined as moles per liter, meaning: \(\sf \large \color{red}{\frac{mol}{L}}\) which means you will be dividing. So, you have 58 grams of NaCl, and divide by the molar mass, this should give you _____ moles of NaCl.

OpenStudy (thadyoung):

Next, divide that number by 4.0 L of H\(_2\)O because molarity is moles of solute divide by liters of solution.